# Effect of noise characterization on the detection of mHz stochastic gravitational waves

### v1

@@ Abstract
Pulsar timing arrays’ hint for a stochastic gravitational-wave background (SGWB) leverages the expectations of a future detection in the millihertz band, particularly with the LISA space mission. However, finding an SGWB with a single orbiting detector is challenging: It calls for cautious modelling of instrumental noise, which is also mainly stochastic. It was shown that agnostic noise reconstruction methods provide robustness in the detection process. We build on previous work to include more realistic instrumental simulations and additional degrees of freedom in the noise inference model and analyze the impact of LISA’s sensitivity to SGWBs. Particularly, we model the two main types of noise sources with separate transfer functions and power spectral density spline fitting. We assess the detectability bounds and their dependence on the flexibility of the noise model and on the prior probability, allowing us to refine previously reported results.
@@ 1 Introduction
The LISA mission is comprised of three-spacecraft forming a triangle constellation, which follows a heliocentric orbit. LISA is scheduled to be launched in the mid-2030s. Inside each spacecraft, two test masses are hosted, being constantly held in free-falling conditions, while their relative distance between the different spacecraft is being monitored via means of laser interferometry [8]. One of the key differences of LISA from ground-based detectors is its interferometric arms cover millions of kilometers and use different laser sources. This requires post-processing the laser link data between the spacecraft in order to subtract the laser frequency noise. This post processing technique is called time-delay interferometry (TDI) [9, 10, 11, 12].
@@
In addition, we have performed our analysis using a full trans-dimensional sampling when using spline models. We compared the outcome of this inference to analytic template models of fixed dimension. We aim to assess the impact of the spectral model choices on the search for SGWB in future LISA data. The instrumental noise of the observatory will not be completely known during mission operations [8]. This is a consequence of the mission requirements, which only set upper bounds on the noise budget. Furthermore, LISA will be a signal-dominated observatory, with multiple types of waveforms populating the data that overlap in time and frequency [8], making it impossible to have GW-quiet data streams. Recent works have also been focusing on the flexible spectral modeling of the noise [14, 16, 17].
Taking into account the above, it is expected that any prior knowledge of instrumental noise will significantly affect our ability to discriminate the potential contributions of individual SGWB to instrumental noise. To investigate this, we compare the performance of searches that use different levels of model flexibility for both signal and noise. To this end, we benchmark the use of analytical versus non-parametric spline models for noise and signal components. We also explore the impact of the prior choice on the inference of the stochastic signal by performing the analysis under both informative and uninformative priors.
@@ 2.1 Frequency-domain data covariance
We showed that for an isotropic background, the power spectral density (PSD) decouples from data correlations so that they can be factored as
@@
We extend the noise model presented in Paper I by considering two different noise sources: the high-frequency optical metrology system (OMS) noise and the low-frequency test mass (TM) noise, each with its own response matrix. If we assume that each noise source is described by an identical PSDs for all interferometers and that the two noise types are completely independent, the noise covariance in Eq.˜2.1 can be decomposed as
@@ 2.2 Noise responses
We start by considering the readout noise in the science interferometric measurement (SCI), which we assimilate in this study to OMS noise, the transformation matrix to form the variables $\eta$ is simply
@@
where $n_{ij,\mathrm{TM}}$ is the projection of the TM jitter onto the sensitive axis and $\mathbf{D}_{ij}$ is the delay operator, which applies a delay $L_{ij}(t)$ equal to the light travel time between spacecraft $i$ and $j$. In the frequency domain, such an operator can be approximated by a complex phasor:
@@
This approximation holds if the delay variations in time can be neglected, for example, when considering orbits with almost right arms, or when considering durations short enough so that the approximation holds. In this work, the former case applies, as we use Keplerian orbits at first order in eccentricity. With our choice of ordering, we have
@@
where we used the short-hand convention $\mathbf{D}_{i_{1}\dots i_{n}}=\Pi_{j=2}^{n-1}\mathbf{D}_{i_{j}i_{j+1}}$. For convenience we rewrite Eq. (2.2) as
@@
We compute the variables $Y_{2}$ and $Z_{2}$ by cyclic permutations of indices. One can rewrite the full TDI transformation in the frequency domain, and express it as a matrix $\mathbf{M}_{\eta\rightarrow d}$:
@@ 3.2 The likelihood model
The appropriate choice for $\nu(f_{j})$ depends mainly on the size of the segment centered in $f_{j}$ and on the window used to taper the time series. We provide details on how to compute it in Ref. [22]. In this work here we have adopted a conservative figure of $\nu(f_{j})=n_{j}/N_{\mathrm{bw}}$, where $N_{\mathrm{bw}}$ is the normalized equivalent noise bandwidth, which depends on the given window function.
@@ 3.3 Model selection
In a Bayesian framework, detecting the presence of a stochastic process can be done through model comparison: one model assumes that the data only contains noise (null hypothesis $H_{0}$), while the other model assumes the presence of a SGWB in addition to the noise (tested hypothesis $H_{1}$). We compare the models by computing their Bayes factor, defined as the ratio of their evidence. The log-Bayes factor is then
@@
where $Z_{i}\equiv\int_{{\boldsymbol{\theta}}_{i}}p\left(\mathbf{\bar{P}}|H_{i}\right)d\boldsymbol{\theta}$ is the evidence of the model under hypothesis $H_{i}$ and $\Theta_{i}$ is the space in which $\boldsymbol{\theta}_{i}$ is allowed to take values. For example, the presence of a SGWB is claimed when the Bayes factor exceeds a given threshold. The appropriate threshold value is usually a challenging quantity to compute. However, it can be chosen following recommendations in the literature (e.g., from [23]), or evaluated by characterizing the Bayes factor distribution from many data realizations, as done in previous works [13, 24]. Other studies can also be utilized in this regard, such as [25, 26].
Running multiple tempered Markov Chain Monte Carlo (MCMC) chains in parallel allows us to use different methods that can approximate the log-evidence $\log Z$ of a given model. The first is the thermodynamic integration [27],
@@
where the expectation $\mathrm{E}_{\beta}$ is taken with respect to the tempered posterior density $p_{\beta}\left(\mathbf{\bar{P}}|\boldsymbol{\theta},H_{i}\right)\propto p\left(\mathbf{\bar{P}}|\boldsymbol{\theta},H_{i}\right)^{\beta}p(\boldsymbol{\theta},H_{i})$. The variable $\beta$ is the inverse temperature of the tempered chain, and $\mathrm{E}_{\beta}$ is the expectation of the chain at temperature $1/\beta$ taken over the parameter space $\Theta$. The second is the so-called stepping stone (SS) algorithm [28], which uses random segments of the MCMC chains to compute the evidence as
$\log\mathcal{Z}=\sum_{k=1}^{N_{T}-1}\log\sum_{i=1}^{n}\log p\left(\mathbf{\bar{P}}|\boldsymbol{\theta},H_{i}\right)^{\beta_{k+1}-\beta_{k}}-(N_{T}-1)\log n,$ (3.6)
where $N_{T}$ is the number of temperatures used in the particular run, and $n$ the number of posterior samples used in the calculation. This method is proven to be more robust in cases where the number of temperatures is relatively small [29].
@@ 3.4 Shape-agnostic spectral models
Following our previous work [13], we can adopt a flexible model for a given PSD. This flexible model is based on spline-interpolation techniques, where the parameters to be estimated are the locations (on the frequency axis) and amplitudes of the spline knots. We choose to use Akima’s splines [34] to represent the log-PSD of each noise single-link component $\alpha$, for which the parameters are the locations and amplitudes of the knot points $x_{i}=\log f_{i},y_{i}=\log S_{i}$. For any log-frequency $x=\log f$ between $x_{i}$ and $x_{i+1}$, we have
@@
The Akima splines exhibit, in general, smoother characteristics than the standard cubic B-spline interpolation functions, which often resulted in spurious PSD estimates for close-by spline knots. In turn, spurious PSD curves may cause numerical issues in the likelihood calculation. The above are reasonable justifications to move away from the spline functions used originally in [13].
One of the common challenges with shape-agnostic models is the actual number and placement of spline knots along the $x$ axis. Depending on the measured data, too many spline knots might result in overfitting issues, while too few might fail to capture the data complexity. As already described in previous subsection 3.3, we solve this problem by adopting trans-dimensional sampling techniques, where the number of knots $k$, their positioning in log-frequency space $x=\log f$, and their amplitudes $\log S_{j,k}$ are determined simultaneously from the data. We remind the reader here that the $j$ index corresponds to the spline number for the given model order $k$.
In [13], the optimal model order for such a configuration was found to be $k=6$. This configuration was suitable for the model assumptions of [13]. In the present work, TM and OMS noises are reflected with different transfer functions (as described in section 3.3), and their interferometric-level PSD $S_{\alpha}$ have distinct shapes. Therefore, a full spectral-shape agnostic model, for the case of equal noise PSDs across interferometers, requires three dynamical spline-interpolation models: two for the overall noise and one for the SGWB signal. To decrease the number of spline knots necessary to characterize the data, we fit deviations from reference log-spectra (taken equal to the injection). This practice restrains the dimensionality of the problem. We have also repeated the analysis on the actual spectra rather than the deviations from the injected ones to verify that we obtain similar results. The difference lies in the much larger required dimensionality of the model, which translates to longer MCMC runs to achieve convergence and therefore an increase of required computational resources.
@@ 3.5.2 Signal template
$S_{\mathrm{GW}}(f)=\Omega_{0}\left(\frac{f}{f_{0}}\right)^{n}\frac{3H_{0}^{2}}{4\pi^{2}f^{3}},$ (3.11)
where the parameter set $\theta_{\mathrm{GW}}\equiv\{\Omega_{0},\,n\}$ contains the amplitude and spectral index of the power law, respectively, $f$ is the frequency, and $H_{0}$ is the Hubble parameter at the present day. The frequency $f_{0}$ is a convention parameter (the so-called pivot frequency) that we set to $f_{0}=0.003~\mathrm{Hz}$.
@@ 3.6 Different priors on the instrumental noise
In this work, we investigate the impact of the noise parameter prior probability distribution on the SGWB search. We compare two types of priors. The first one is uninformative: it is a wide uniform distribution centered on the true injected noise spectrum, which is chosen as a baseline for comparison with our earlier work [13]. This means that, for the case of the flexible shape-agnostic spline model, the knot amplitudes are allowed only within the bounds of $\log_{10}S_{j,k}\sim\mathcal{U}[-1,\,1]$. We also repeat the analysis with an informative prior, which is based on the skewed normal distribution [35], again centered around the true injected noise spectra. The skewed normal distribution for a random variable $x$ is defined as
@@ 3.7 Priors on the stochastic gravitational wave signal
As already described, we adopt two types of models for the signal analysis: a template-based and a spline-based description.
When using the template model described in 3.5.2, the priors for the SGWB signal are rather wide uninformative priors of $\log\Omega_{0}\sim\mathcal{U}[-20,\,-9]$ and $n\sim\mathcal{U}[-5,\,5]$, in order to cover the complete parameter space which potentially yields detectable signals.
When the agnostic spline model described in 3.4 is used to fit the signal, we choose a wide uniform prior with a maximum allowed power as $\log S_{\mathrm{GW}}(f,\boldsymbol{\theta}_{\mathrm{GW}})\sim\mathcal{U}[-135,\,-71]$ for all frequencies.
We expect improved detectability capabilities for a given SGWB signal when the prior assumptions are informative and when the fitting models are templates depending on fewer parameters. We verify this expectation in the sections below.
@@ 4 Simulations and analysis
As in [13], we choose to work with the expectation values for the data instead of realizations. In particular, if $\mathbf{\bar{Y}}=\operatorname{E}_{\theta^{\star}}[\mathbf{Y}]$ is the expectation of the data $\mathbf{Y}$ under the true hypothesis, then $\overline{\log\mathcal{B}_{10}}=\mathcal{B}_{\mathrm{thresh}}$ yields a conservative criterion for detection, for a given $\mathcal{B}_{\mathrm{thresh}}$. We also choose the $\mathcal{B}_{\mathrm{thresh}}=100$ as the threshold for confident signal detection.
Figure 2: Bayes Factors $B_{10}$ between the noise+signal $H_{1}$ and the noise-only $H_{0}$ hypotheses, for different injections of SGWB, under the assumptions of various priors on the noise model, and also different model types (parametric versus non-parametric). The first row contains the results of the runs where non-parametric models were used for all the noise and signal models, while the second row shows the results where a parametric model was used only for the signal. The last row shows the results obtained when using parametric models for all contributions to the covariance matrix. The first (respectively the second) column shows the results of the runs where the informative $\mathcal{SN}$ (respectively non-informative $\mathcal{U}$) prior was used on the instrumental noise component.
@@
As a variety of stochastic signals from cosmological or astrophysical origins are predicted to follow a power-law at the $\mathrm{mHz}$ range  [39], we adopt this type of model for our injections. Thus, we create a grid of amplitudes and spectral indices $\theta_{\mathrm{GW}}\equiv\{\Omega_{0},\,n\}$, essentially covering the parameter space that is potentially covered by the LISA sensitivity. The grid of injections spans between $10^{-15}<\Omega_{0}<5\times 10^{-12}$ and 3<n<3-3<n<3. For each pair of $\theta_{\mathrm{GW}}$, we run six types of inferences: parametric (analytic) and non-parametric (spline) models for signal and noise, both with informative and uninformative prior assumptions. For each injection, we also run the noise-only case, intending to compute the Bayes factors and assess the detectability of the given signal under the different modelling assumptions. Injections always contain noise and signal.
Finally, as already mentioned, when using the non-parametric spline approach, we always infer the optimal number of spline knots using RJ sampling methods. In the end, for each run we also compute the evidence for the given model using the Stepping Stone algorithm, as introduced in section 3.3. The evidence computation of the full trans-dimensional problem is achievable because only the temperatures and likelihoods for each RJ-MCMC walker are needed [31]. The noise-only versus the noise-plus-signal hypotheses are then tested by computing the Bayes Factor as the ratio of the evidences calculated. The sampler was configured to run MCMC chains for $50$ walkers, for each of the $30$ temperatures used in the adaptive parallel tempering scheme. After a short burn-in period, during sampling computations each second MCMC point was kept in the chain until a total $10^{4}$ MCMC samples was gathered for each walker. The results of our analysis are summarized in figure 2 and in the sections below. Example reconstructions of the recovered signals can be seen in figure 4.

### v2

@@ Abstract
Pulsar timing arrays’ hint for a stochastic gravitational-wave background (SGWB) leverages the expectations of a future detection in the millihertz band, particularly with the LISA space mission. However, finding a SGWB with a single orbiting detector is challenging: It calls for cautious modelling of instrumental noise, which is also mainly stochastic. It was shown that agnostic noise reconstruction methods provide robustness in the detection process. We build on previous work to include more realistic instrumental simulations and additional degrees of freedom in the noise inference model and analyze the impact of LISA’s sensitivity to SGWBs. Particularly, we model the two main types of noise sources with separate transfer functions and power spectral density spline fitting. We assess the detectability bounds and their dependence on the flexibility of the noise model and on the prior probability, allowing us to refine previously reported results.
@@ 1 Introduction
The LISA mission is comprised of three spacecraft forming a triangle constellation, which follows a heliocentric orbit. LISA is scheduled to be launched in the mid-2030s. Inside each spacecraft, two test masses are hosted, being constantly held in free-falling conditions, while their relative distance between the different spacecraft is being monitored via means of laser interferometry [8]. One of the key differences of LISA from ground-based detectors is its interferometric arms cover millions of kilometers and use different laser sources. This requires to post-process the laser link data between the spacecraft in order to subtract the laser frequency noise. This post-processing technique is called time-delay interferometry (TDI) [9, 10, 11, 12].
@@
In addition, we have performed our analysis using a full trans-dimensional sampling when using spline models, a technique which was first introduced in the context of GW data analysis in [16]. We compared the outcome of this inference to analytic template models of fixed dimension. We aim to assess the impact of the spectral model choices on the search for SGWB in future LISA data. The instrumental noise of the observatory will not be completely known during mission operations [8]. This is a consequence of the mission requirements, which only set upper bounds on the noise budget. Furthermore, LISA will be a signal-dominated observatory, with multiple types of waveforms populating the data that overlap in time and frequency [8], making it impossible to have GW-quiet data streams. Recent works have also been focusing on the flexible spectral modeling of the noise [14, 17, 18].
Taking into account the above, it is expected that any prior knowledge of instrumental noise will significantly affect our ability to discriminate the potential contributions of individual SGWB to instrumental noise. To investigate this, we compare the performance of searches that use different levels of model flexibility for both signal and noise. To this end, we benchmark the use of analytical versus flexible spline models for noise and signal components. We also explore the impact of the prior choice on the inference of the stochastic signal by performing the analysis under both informative and uninformative priors.
@@ 2.1 Frequency-domain data covariance
We showed that for an isotropic background, the power spectral density (PSD) can be factored o as
@@
We extend the noise model presented in Paper I by considering two different noise sources: the high-frequency optical metrology system (OMS) noise and the low-frequency test mass (TM) noise, each with its own response matrix. If we assume that each noise source is described by an identical PSD for all interferometers and that the two noise types are completely independent, the noise covariance in Eq.˜2.1 can be decomposed as
@@ 2.2 Noise responses
We start by considering the readout noise in the science interferometric measurement (SCI), which we assimilate in this study to OMS noise. The transformation matrix to form the variables $\eta$ is simply
@@
where $n_{ij,\mathrm{TM}}$ is the projection of the TM jitter onto the sensitive axis and $\mathbf{D}_{ij}$ is the delay operator, which applies a delay $L_{ij}(t)$ equal to the light travel time from spacecraft $j$ to $i$. In the frequency domain, such an operator can be approximated by a complex phasor:
@@
This approximation holds if the delay variations in time can be neglected, for example, when considering orbits with constant arms, or when considering durations short enough so that the approximation holds. In this work, the former case applies, as we use Keplerian orbits at first order in eccentricity. With our choice of ordering, we have
@@
where we used the short-hand convention $\mathbf{D}_{i_{1}\dots i_{n}}=\Pi_{j=1}^{n-1}\mathbf{D}_{i_{j}i_{j+1}}$. For convenience we rewrite eq. (2.2) as
@@
We compute the variables $Y_{2}$ and $Z_{2}$ by cyclic permutations of indices ($1\rightarrow 2\rightarrow 3\rightarrow 1)$. One can rewrite the full TDI transformation in the frequency domain, and express it as a matrix $\mathbf{M}_{\eta\rightarrow d}$:
@@ 3.2 The likelihood model
The appropriate choice for $\nu(f_{j})$ depends mainly on the size of the segment centered in $f_{j}$ and on the window used to taper the time series. We provide details on how to compute it in Ref. [23]. In this work here we have adopted a conservative figure of $\nu(f_{j})=n_{j}/N_{\mathrm{bw}}$, where $N_{\mathrm{bw}}$ is the normalized equivalent noise bandwidth, which depends on the given window function [24].
@@ 3.3 Model selection
In a Bayesian framework, detecting the presence of a stochastic process can be done through model comparison: one model assumes that the data only contains noise (null hypothesis $H_{0}$), while the other model assumes the presence of a SGWB in addition to the noise (tested hypothesis $H_{1}$). We compare the models by computing their Bayes factor, defined as the ratio of their evidences. The log-Bayes factor is then
@@
Running multiple tempered Markov Chain Monte Carlo (MCMC) chains in parallel allows us to use different methods that can approximate the log-evidence $\log Z_{i}$ of a given model. The first is the thermodynamic integration [29],
[…]
where the expectation $\mathrm{E}_{\beta}$ is taken with respect to the tempered posterior density $p_{\beta}\left(\mathbf{\bar{P}}|\boldsymbol{\theta}_{i},H_{i}\right)\propto p\left(\mathbf{\bar{P}}|\boldsymbol{\theta}_{i},H_{i}\right)^{\beta}p(\boldsymbol{\theta}_{i},H_{i})$. The variable $\beta$ is the inverse temperature of the tempered chain, and $\mathrm{E}_{\beta}$ is the expectation of the chain at temperature $1/\beta$ taken over the parameter space $\Theta_{i}$. The second is the so-called stepping stone (SS) algorithm [30, 31], which uses random segments of the MCMC chains to compute the evidence as
$\log Z_{i}=\sum_{k=1}^{N_{T}-1}\log\sum_{i=1}^{n}\log p\left(\mathbf{\bar{P}}|\boldsymbol{\theta}_{i},H_{i}\right)^{\beta_{k+1}-\beta_{k}}-(N_{T}-1)\log n,$ (3.6)
where $N_{T}$ is the number of temperatures used in the particular run, and $n$ the number of posterior samples used in the calculation. This method is proven to be more robust in cases where the number of temperatures is relatively small [32]. Recently there have been more methods been introduced in the literature, such as the one of [33], but we choose to work with the SS algorithm, due to its availability within the Eryn package.
@@ 3.4 Shape-agnostic spectral models
Following our previous work [13], we can adopt a flexible model for a given PSD. This flexible model is based on spline-interpolation techniques, where the parameters to be estimated are the locations (on the frequency axis) and amplitudes of the spline knots. We choose to use Akima splines [38] to represent the log-PSD of each noise component $\alpha$, for which the parameters are the locations and amplitudes of the knot points $x_{i}=\log f_{i},y_{i}=\log S_{i}$. For any log-frequency $x=\log f$ between $x_{i}$ and $x_{i+1}$, we have
@@
The Akima splines are considered to be very stable locally, and therefore do not exhibit Gibbs-like ringing or overshoot, in contrast to the standard cubic B-spline interpolation functions, which often resulted in spurious PSD estimates for close-by spline knots. In turn, spurious PSD curves may cause numerical issues in the likelihood calculation. The above are reasonable justifications to move away from the spline functions used originally in [13].
One of the common challenges with shape-agnostic models is the actual number and placement of spline knots along the $x$ axis. Depending on the measured data, too many spline knots might result in overfitting issues, while too few might fail to capture the data complexity. As already described in previous subsection 3.3, we solve this problem by adopting trans-dimensional sampling techniques, where the number of knots $k$, their positioning in log-frequency space $x_{j,k}=\log f_{j,k}$, and their amplitudes $\log S_{j,k}$ are determined simultaneously from the data. We remind the reader here that the $j$ index corresponds to the spline number for the given model order $k$.
In [13], the optimal model order for such a configuration was found to be $k=6$. This configuration was suitable for the model assumptions of [13]. In the present work, TM and OMS noises are reflected with different transfer functions (as described in section 3.3), and their interferometric-level PSD $S_{\alpha}$ have distinct shapes. Therefore, a full spectral-shape agnostic model, for the case of equal noise PSDs across interferometers, requires three dynamical spline-interpolation models: two for the overall instrument noise and one for the SGWB signal. To decrease the number of spline knots necessary to characterize the data, we fit deviations from reference log-spectra (taken equal to the injection). This practice restrains the dimensionality of the problem. We have also repeated the analysis on the actual spectra rather than the deviations from the injected ones to verify that we obtain similar results. The difference lies in the much larger required dimensionality of the model, which translates to longer MCMC runs to achieve convergence and therefore an increase of required computational resources.
@@ 3.5.2 Signal template
$S_{\mathrm{GW}}(f)=\Omega_{\mathrm{GW}}\left(\frac{f}{f_{0}}\right)^{n}\frac{3H_{0}^{2}}{4\pi^{2}f^{3}},$ (3.11)
where the parameter set $\boldsymbol{\theta}_{\mathrm{GW}}\equiv\{\Omega_{\mathrm{GW}},\,n\}$ contains the amplitude and spectral index of the power law, respectively, $f$ is the frequency, and $H_{0}$ is the Hubble parameter at the present day. The frequency $f_{0}$ is a convention parameter (the so-called pivot frequency) that we set to $f_{0}=0.003~\mathrm{Hz}$.
@@ 3.6 Different priors on the instrumental noise
In this work, we investigate the impact of the noise parameter prior probability distributions on the SGWB search. We compare two types of priors. The first one is uninformative: it is a wide uniform distribution centered on the true injected noise spectrum, which is chosen as a baseline for comparison with our earlier work [13]. This means that, for the case of the flexible shape-agnostic spline model, the knot amplitudes are allowed only within the bounds of $\log_{10}S_{j,k}\sim\mathcal{U}[-1,\,1]$. We also repeat the analysis with an informative prior, which is based on the skewed normal distribution [39], again centered around the true injected noise spectra. The skewed normal distribution for a random variable $x$ is defined as
@@ 3.7 Priors on the stochastic gravitational wave signal
Figure 2: Bayes Factors $\mathcal{B}_{10}$ between the noise-and-signal $H_{1}$ and the noise-only $H_{0}$ hypotheses, for different injections of SGWB, under the assumptions of various priors on the noise model, and also different model types (parametric versus flexible). The first row contains the results of the runs where flexible models were used for all the noise and signal models, while the second row shows the results where a parametric model was used only for the signal. The last row shows the results obtained when using parametric models for all contributions to the covariance matrix. The first (respectively the second) column shows the results of the runs where the informative $\mathcal{SN}$ (respectively non-informative $\mathcal{U}$) prior was used on the instrumental noise component.
As already described, we adopt two types of models for the signal analysis: a template-based and a spline-based description. When using the template model described in 3.5.2, the priors for the SGWB signal are rather wide uninformative priors of $\log\Omega_{\mathrm{GW}}\sim\mathcal{U}[-20,\,-9]$ and $n\sim\mathcal{U}[-5,\,5]$, in order to cover the complete parameter space that potentially yields detectable signals. When the agnostic spline model (described in 3.4) is used to fit the signal, we choose a wide uniform prior for the spline knots, with a maximum allowed power as $\log S_{\mathrm{GW}}(f,\boldsymbol{\theta}_{\mathrm{GW}})\sim\mathcal{U}[-135,\,-71]$ for all frequencies. We expect improved detectability capabilities for a given SGWB signal when the prior assumptions are informative and when the fitting models are templates depending on fewer parameters. We verify this expectation in the sections below.
@@ 4 Simulations and analysis
As in [13], we choose to work with the expectation values for the data instead of realizations. In particular, if $\mathbf{\bar{Y}}=\operatorname{E}_{\boldsymbol{\theta}^{\star}}[\mathbf{Y}]$ is the expectation of the data $\mathbf{Y}$ under the true hypothesis, then $\overline{\log\mathcal{B}_{10}}=\mathcal{B}_{\mathrm{thresh}}$ yields a conservative criterion for detection, for a given $\mathcal{B}_{\mathrm{thresh}}$. We also choose $\mathcal{B}_{\mathrm{thresh}}=100$ as the threshold for confident signal detection.
@@
As a variety of stochastic signals from cosmological or astrophysical origins are predicted to follow a power-law at the $\mathrm{mHz}$ range  [43], we adopt this type of model for our injections. Thus, we create a grid of amplitudes and spectral indices $\boldsymbol{\theta}_{\mathrm{GW}}\equiv\{\Omega_{\mathrm{GW}},\,n\}$, essentially covering the parameter space that is potentially covered by the LISA sensitivity. The grid of injections spans between $10^{-15}<\Omega_{\mathrm{GW}}<5\times 10^{-12}$ and 3<n<3-3<n<3. For each pair of $\boldsymbol{\theta}_{\mathrm{GW}}$, we run six types of inferences: parametric (analytic) and flexible (spline) models for signal and noise, both with informative and uninformative prior assumptions. For each injection, we also run the noise-only case, intending to compute the Bayes factors and assess the detectability of the given signal under the different modelling assumptions. Injections always contain noise and signal.
Finally, as already mentioned, when using the flexible spline approach, we always infer the optimal number of spline knots using RJ sampling methods. In the end, for each run we also compute the evidence for the given model using the Stepping Stone algorithm, as introduced in section 3.3. The evidence computation of the full trans-dimensional problem is achievable because only the temperatures and likelihoods for each RJ-MCMC walker are needed [35]. The noise-only versus the noise-plus-signal hypotheses are then tested by computing the Bayes Factor as the ratio of the evidences calculated. The sampler was configured to run MCMC chains for $50$ walkers, for each of the $30$ temperatures used in the adaptive parallel tempering scheme. After a short burn-in period, during sampling computations each second MCMC point was kept in the chain until a total $10^{4}$ MCMC samples was gathered for each walker. The results of our analysis are summarized in figure 2 and in the sections below. Example reconstructions of the recovered signals can be seen in figure 4.